OpenStudy (anonymous):

I need the steps to this equation.? I need the steps to this equation.? A person invested $6400 for 1 year, part at 5%, part at 10%, and the remainder at 13% was $687. The amount of money invested at 13% was $400 more than the amounts invested at 5% and 10% combined. Find the amount invested at each rate.

OpenStudy (pfenn1):

Let x = $ invested at 5% y = $ invested at 10% 686 = $ invested at 13% We know that 6400 was invested so \[x+y+686=6400\]We also know that $687 is $400 more than the amounts invested at 5% and 10% combined, or\[687-400=x+y\]So now we have two equations and two unknowns. Can you solve this?